# Improved Patch Denoising Diffusion Probabilistic Models for Magnetic Resonance Fingerprinting

### v1

@@ 4.9 Estimation of missing timeframes
Table 3: Reconstruction errors for TSMI and T1/T2 parameter maps using IDDPM-MRF (ours) and DRUNet models trained on two tasks: 1) TSMI restoration only, and 2) TSMI restoration and mapping a longer timeframe.
We observed MRF-IDDPM offered a marginal gain in the reconstruction metrics in $l=1000$ (gain of $\sim 1\%$ in T1/T2 MAPE and $\sim 7\%$ in NRMSE for TSMI). This could be because longer timeseries give more discriminant signals for different T1/T2 values and so reconstructions were able to slightly benefit from this. However we did not observe such gain in the standard CNN implementation, where DRUNet performs better at reconstructing TSMIs at $l=200$, as well as the resulting Q-Maps, but its performance is weaker at $l=1000$. MRF-IDDPM outperforms DRUNet in both tasks.
4.10 Uncertainty Maps
Conditional diffusion models are shown to be useful to additionally produce a notion of uncertainty map alongside their reconstructed images (for example, see [45, 35]). This uncertainty can be visualized as a variance map of the output images obtained from multiple runs of the reconstruction (sampling the reverse diffusion) process in Algorithm 2. Each run starts with a different random draw of the initial noise $\mathbf{x}_{T}\sim\mathcal{N}(\textbf{0},\textbf{Id})$, but all use the same conditioning image $\mathbf{x}_{c}$. The hypothesis is that, in a robust and well-trained model, regions with high variance will correlate with areas of large absolute errors in the reconstruction. Moreover, averaging reconstructions across multiple samples is expected to reduce error metrics. To test this, we ran the reverse process to reconstruct 10 TSMIs (and the corresponding T1 and T2 maps), using the same low-quality conditioning image but starting with different initial points $\mathbf{x}_{T}$ each time. We then computed the pixel-wise standard deviation across the 10 sampled outputs and compared it to the absolute error between the averaged reconstructions and the reference. Figure 5 shows the averaged T1 and T2 reconstructions, along with the variance and absolute error maps for the selected brain slice in Figure 1. As suggested by the hypothesis, the figures demonstrate clear correlations between the variance and error maps, with high errors occurring primarily in the cerebrospinal fluid and surrounding fat/skull areas, and less so in white and gray matter.
Fig. 5: Average, variance, and absolute error maps of T1 (left), T2 (right) across 10 samples of reverse diffusion process (i.e. 10 reconstruction runs each with a different initializing point $\mathbf{x}_{T}$ but all with same condition $\mathbf{x}_{c}$). The variance maps, obtained as described in [35], display pixel-wise standard deviations calculated from the 10 samples. The absolute error map represents the absolute difference between the averaged map (left) and the reference (Fig. 1).
@@ 4.11 Limitations and Future Work
Our study was restricted to a small in-house dataset of healthy subjects, and thus generalization to other datasets is yet to be examined, including data containing pathologies, brain diseases, or different anatomical regions than brain. Furthermore, our approach was tested on 2D imaging data and considering future extensions to 3D MRF acquisitions e.g. [4, 5, 46] would require accounting for spatial context and correlations across 3D volumes, as well as increased computational resources. Processing images at the patch level as proposed, or using latent space approaches like encoder-decoder models could help manage the high dimensionality and memory demands for 3D reconstructions. Lastly, in our current pipeline, Bloch response constraints are applied during the parameter quantification stage, after TSMI reconstruction. While achieving good performance, further improvements might be achieved by incorporating these constraints, along with k-space consistency, during the TSMI reconstruction process itself (see e.g. [17, 18, 47] for non-DDPM based MRF models that account for these constraints, also [48, 49] for MRI reconstructions DDPM models that account for k-space consistency).
@@ 5 Conclusions
This paper introduces a novel IDDPM-based model for the enhanced reconstruction of parameter maps from fast transient-state qMRI techniques, such as MRF. We evaluated our method on in-vivo brain scans from healthy volunteers, where the MRF acquisition sequence (FISP) was retrospectively truncated by a factor of 5, i.e. 5-fold acceleration. Both qualitative and quantitative results demonstrated that the proposed approach, IDDPM-MRF, outperforms established deep learning and compressed sensing algorithms for MRF reconstruction. To reduce computational overhead, we employed an efficient DDPM sampling strategy and utilized patch-level image processing during both training and reconstruction, allowing these steps to be performed on modest GPUs within a reasonable runtime.

### v2

@@ 4.5.4 Estimation of missing time frames
Table 3: Reconstruction errors for TSMI and T1/T2 parameter maps using MRF-IDDPM and DRUNet models trained on two tasks: 1) TSMI restoration only and 2) TSMI restoration and mapping to a longer timeframe.
The MRF-IDDPM demonstrates a marginal improvement in reconstruction metrics for Task 2, with gains of $\sim 1\%$ in T1/T2 MAPE and $\sim 7\%$ in TSMI NRMSE. This could be because longer time series give more discriminant signals for different T1/T2 values and so reconstructions were able to benefit from this. However, this advantage was not observed in the standard CNN implementation. DRUNet performed better in reconstructing TSMIs at $l=200$, but its performance declined on Task 2. Notably, MRF-IDDPM outperformed DRUNet across both tasks.
4.5.5 Uncertainty Maps
Conditional diffusion models are shown to be useful to additionally produce a notion of uncertainty map alongside their reconstructed images (for example, see [52, 40]). This uncertainty can be visualized as a variance map of the output images obtained from multiple runs/samples of the reconstruction process in Algorithm 2. Each run starts with a different random draw of the initial noise $\mathbf{x}_{T}\sim\mathcal{N}(\textbf{0},\textbf{Id})$, but all use the same conditioning image $\mathbf{x}_{c}$. This allows for the inspection of the model by assessing the variations across the reconstructed samples. To show this, we ran the reverse diffusion process to reconstruct 10 TSMIs (and the corresponding T1 and T2 maps), using the same low-quality conditioning image but starting with different initial random noise $\mathbf{x}_{T}$ each time. We then computed the pixel-wise standard deviation across the 10 sampled outputs and compared it to the absolute error between the averaged reconstructions and the reference. Figure 5 shows the averaged T1 and T2 reconstructions, along with the variance and absolute error maps for three selected brain slices in Figure 1. The figures demonstrate clear correlations between the variance and error maps, with higher errors occurring in regions of higher variance. Large errors are primarily occurring in the cerebrospinal fluid and surrounding fat/skull areas, and less so in white and gray matter.
@@ 4.6 Limitations and Future Work
Our study was restricted to a small in-house dataset of healthy subjects, and thus generalization to other datasets is yet to be examined, including data containing pathologies, brain diseases, or different anatomical regions than the brain. Furthermore, our approach was tested on 2D imaging data and considering future extensions to 3D MRF acquisitions e.g. [4, 5, 53] would require accounting for spatial context and correlations across 3D volumes, as well as increased computational resources. Processing images at the patch level as proposed, or using latent space approaches like encoder-decoder models could help manage the high dimensionality and memory demands for 3D reconstructions. Lastly, in our current pipeline, Bloch response constraints were applied during the parameter quantification stage, after TSMI reconstruction. While achieving good performance, further improvements might be achieved by incorporating these constraints, along with k-space consistency, during the TSMI reconstruction process itself (see e.g. [22, 23, 54] for non-DDPM based MRF models that account for these constraints, also [55, 56] for MRI reconstructions DDPM models that account for k-space consistency).
@@ 5 Conclusions
This paper introduced a novel IDDPM-based model for the enhanced reconstruction of parameter maps from fast transient-state qMRI techniques, such as MRF. We evaluated our method on in-vivo brain scans from healthy volunteers, where the MRF acquisition sequence was retrospectively truncated by a factor of 5, i.e. 5-fold acceleration. Both qualitative and quantitative results demonstrated that the proposed approach, MRF-IDDPM, outperforms established deep learning and conventional compressed sensing algorithms for MRF reconstruction. The MRF-IDDPM’s probabilistic framework allows for sampling variations across reconstructed images, which can be useful for generating uncertainty maps and to assess confidence in the reconstruction at each pixel. To reduce computational overhead, we employed an efficient DDPM sampling strategy and utilized patch-level image processing during both training and reconstruction, allowing these steps to be performed on modest GPUs within a reasonable runtime.